GEOMETRIC ZETA FUNCTIONS, L^{2}-THEORY,

AND COMPACT SHIMURA MANIFOLDS

Anton Deitmar

Abstract. We define geometric zeta functions for locally symmetric spaces as generalizations of the zeta functions of Ruelle and Selberg. As a special value at zero we obtain the Reidemeister torsion of the manifold. For hermitian spaces these zeta functions have as special value the quotient of the holomorphic torsion of Ray and Singer and the holomorphic L^{2}-torsion, where the latter is defined via the L^{2}-theory of Atiyah. For higher fundamental rank twisted torsion numbers appear.

Typeset by AMS-TEX